Week of 190527

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Alberto (Monitoring), Borja (Chair, Monitoring), Dmitry (LHCb), Ivan (ATLAS), Julia (WLCG), Maarten (ALICE), Olga (Computing), Vincent (Security), Vladimir (LHCb), ?? (Storage)
  • remote: Christoph (CMS), Di (TRIUMF), Dmytro (NDGF), Onno (NL-T1), Pepe (PIC), Xavier (KIT)

Experiments round table:

  • ATLAS reports ( raw view) -
    • VOMS
      • Not available.
        • Last Wednesday for several hours proxy initiation / renewal was not working.
        • Error message: "None of the contacted servers for atlas were capable of returning a valid AC for the user"
        • No explanation what happened.(?)
        • Status: Solved.
      • Overload.
        • Followed at INC:1993132
        • Experts: Service is overloaded with requests by:
          • Unknown IPs. No one is looking into those (?)
          • ATLAS Panda. Under investigation.
          • Weizmann site. Discovered by ATLAS (?). Ticket submitted by ATLAS (?)
        • Status: Ongoing understanding attempt
        • Nothing changed on our side
      • (Points to be addressed by the VOMS team denoted with (?))
    • Tier-1s FYI: ATLAS Tape deletion campaign was finished.
    • A gentle reminder of the CentOS7 deployment deadline - 01.06.2019. Please update your site status here.

Maarten reported he has been involved with the VOMS investigation, there was and increase on the load leading to a reach on the limit of sessions, as a workaround the limit has been increased from 200 to 300 and the investigation is ongoing since this was the first time this was seen.

Ticket will continue investigating the issues.

  • CMS reports ( raw view) -
    • CPU utilization ~210k cores
      • Less than usual mainly, because HLT farm is off
    • Otherwise no major issues

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Smooth running at ~100K jobs, Usual activity
      • User jobs, MC productions, and WG productions this week
    • issues which are not significant, but potentially may be of interest to other experiments:
      • in progress: Poor transfer efficiency from CERN WN to outside storage GGUS:141112
      • Users getting <TNetXNGFile::Open>: [FATAL] Auth failed at RAL GGUS:141262

Maarten asked whether this is a CEPH related issue, Vladimir said it seems to be some GRID map file not being updated in RAL nodes.

Sites / Services round table:

  • ASGC: NC
  • BNL: NTR
  • CNAF: NC
  • EGI: NC
  • FNAL: NC
  • IN2P3: Site will be on scheduled maintenance next Tuesday, June 4th: dCache upgrade to 4.2, security update on WN/ Scheduled downtime on CEs and SEs for the whole day.
  • JINR: NTR
  • KISTI: NC
  • KIT: We plan to have half of our batch farm nodes migrated to SL7 by Wednesday this week.

Ivan asked what was the plan for the rest, migration will continue at the pace of a couple of racks per day, estimated finishing time will be within June.

  • NDGF: NTR
  • NL-T1: NTR
  • NRC-KI: NC
  • OSG: NC
  • PIC: Planned downtime for tomorrow, from 28-05-2019 08:00 (PIC local time) until 28-05-2019 17:00 (PIC local time). dCache will be upgraded to 4.2.32-1 and HTCondor to 8.8.2.
  • RAL: NC
  • TRIUMF: NTR

  • CERN computing services: NTR
  • CERN storage services: NTR
  • CERN databases: NC
  • GGUS:
    • A new release was applied last Wed, along with the usual test alarms
  • Monitoring: NTR
  • MW Officer: NC
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r20 < r19 < r18 < r17 < r16 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r20 - 2019-05-27 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback